OpenStudy (anonymous):

if the speed o an object in uniform circular motion is doubled and the radius remains constant then the centripetal acceleration increases by a factor of

OpenStudy (anonymous):

\[F _{c} = mv ^{2}/r\]

OpenStudy (anonymous):

if speed is doubled means 2V then \[F = m (2V)^{2}/r \]

OpenStudy (anonymous):

\[F = 4mv ^{2}/r\]

OpenStudy (anonymous):

F = ma = 4mv^2/r So acceleration increases by a factor of 4 .

OpenStudy (anonymous):

@Fufii

OpenStudy (anonymous):

a=v2/r double v = quadruple a